What is the charge of 96 amu of S^2- ?
1
Expert's answer
2020-11-12T06:36:08-0500

Atomic mass of sulfide ion = 32amu

Number of sulphide ions of 96 amu=3 ions

Charge on 1 sulphide ion= -2

Charge on 3 sulphide ion =3×(-2) = -6

Charge on 6 electrons = 6×1.6×10-19C

= 9.6×10-19C